# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v1

@@ 4 When will Personalization Improve Statistical Accuracy
The loss function $\ell(\cdot,z)$ is $\mu$-strongly convex for each local sample average, i.e. $\sum_{j\in[n_{i}]}\ell(x,z_{ij})/n_{i}\geq\sum_{j\in[n_{i}]}\ell(y,z_{ij})/n_{i}+\langle\sum_{j\in[n_{i}]}\nabla\ell(y,z_{ij})/n_{i},y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d},i\in[m].$
Assumption 3 (Bound Gradient Variance at Optimum).
@@
Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis. To facilitate the analysis, we first define some notations. Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models. A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$. In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy. Here the expectation is taken with respect to all data. Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error. They are equivalent to our metrics under the assumptions used in this paper. We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error. See Appendix A.2 for the proof.
@@ 5 When will Personalization Improve Communication Efficiency
Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in Appendix B.2.
@@ 6 Simulation
In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions. We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3). Detailed descriptions of the experimental setup are available in Section 6.1, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
@@ 6.2 Statistical Aspect: Balancing Local and Global Learning
We now explore how FedProx balances between GlobalTrain and LocalTrain by achieving different levels of personalization. For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx LABEL:sec:additional_exp. We run all methods to a stable convergence and measure the statistical error. The top row of Figure 1 shows the performance of these methods under different impact of statistical heterogeneity.

### v2

@@ 4 When will Personalization Improve Statistical Accuracy
Each loss function $L_{i}$ is $\mu$-strongly convex, i.e., $L_{i}({\bm{w}},S_{i})\geq L_{i}({\bm{w}}^{\prime},S_{i})+\langle\nabla L_{i}({\bm{w}}^{\prime},S_{i}),{\bm{w}}-{\bm{w}}^{\prime}\rangle+\frac{\mu}{2}\|{\bm{w}}-{\bm{w}}^{\prime}\|^{2},\quad\forall{\bm{w}},{\bm{w}}^{\prime}\in\mathbb{R}^{d},i\in[m].$
Assumption 3 (Bound Gradient Variance at the Optimum).
@@
Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis. To facilitate the analysis, we first define some notations. Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models. A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$. In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy. Here the expectation is taken with respect to all data. Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error. They are equivalent to our metrics under the assumptions used in this paper. We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error. See A.2 for the proof.
@@ 5 When will Personalization Improve Communication Efficiency
Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in B.1.
@@ 6 Simulation
In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions. We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3). Detailed descriptions of the experimental setup are available in Section 6.1.
@@ 6.2 Statistical Aspect: Balancing Local and Global Learning
We now explore how FedProx balances between GlobalTrain and LocalTrain by achieving different levels of personalization. For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx C. We run all methods to a stable convergence and measure the statistical error. The top row of Figure 1 shows the performance of these methods under different impact of statistical heterogeneity.